# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ Abstract.
Let kq denote the very effective cover of the motivic Hermitian K-theory spectrum. We analyze the ring of cooperations $\pi^{\mathbb{R}}_{**}(\textup{kq}\otimes\textup{kq})$ in the stable motivic homotopy category $\text{SH}(\mathbb{R})$, giving a full description in terms of Brown–Gitler comodules. To do this, we decompose the $E_{2}$-page of the motivic Adams spectral sequence and show that it must collapse. The description of the $E_{2}$-page is accomplished by a series of algebraic Atiyah–Hirzebruch spectral sequences which converge to the summands of the $E_{2}$-page. Along the way, we prove a splitting result for the very effective symplectic K-theory ksp over any base field of characteristic not two.
@@ 1.1. Motivation
In stable homotopy theory, one of the most powerful tools we have for computing the homotopy groups of spheres is the Adams spectral sequence. For a ring spectrum E, the E-Adams spectral sequence (E-ASS) has signature
$E_{1}^{s,f}=\pi_{s+f}(\text{E}\otimes\overline{\text{E}}^{\otimes f})\implies\pi_{s}(\mathbb{S}_{\text{E}}^{\wedge}),$
where $\overline{\text{E}}$ is the cofiber of the unit map $\mathbb{S}\to\text{E}$. Under mild assumptions on E, this spectral sequence strongly converges to the homotopy groups of $\mathbb{S}_{\text{E}}^{\wedge}$, the E-nilpotent completion of the sphere spectrum [Rav86]. To begin to compute with the E-ASS, one must understand the smash powers $\text{E}^{\otimes f}$.
@@
The spectra $\textup{H}\mathbb{Z}_{k}^{cl}$ are finite complexes known as the integral Brown–Gitler spectra [BG73, GJM86].
In stable motivic homotopy theory, a primary object of investigation is the homotopy ring of the motivic sphere spectrum. This ring has rich connections with the Grothendieck-Witt ring of symmetric bilinear forms [Mor12], Milnor K-theory [BBX25], and Hermitian K-theory [RSØ19, RSØ24]. Similar to the classical case, one of the most powerful and well-studied tools is the motivic Adams spectral sequence. For a motivic ring spectrum E, there is an E-motivic Adams spectral sequence (E-mASS) which can be used to compute these homotopy groups that, under mild conditions, converges strongly to the homotopy groups of the E-nilpotent completion $\mathbb{S}_{\text{E}}^{\wedge}$ [DI10, HKO11]. This takes the form
$E_{1}^{s,f,w}=\pi^{F}_{s+f,w}(\textup{E}\otimes\overline{\textup{E}}^{\otimes f})\implies\pi_{s,w}^{F}(\mathbb{S}_{\textup{E}}^{\wedge}).$
To begin to compute with the E-mASS, one must understand the smash powers $\textup{E}^{\otimes f}$.
Let kq denote the very effective cover of the Hermitian K-theory spectrum KQ [ARØ20] (see Section 3 for details on the very effective slice filtration). In [CQ21], Culver–Quigley analyze the kq-mASS at the prime 2 in $\mathbb{C}$-motivic homotopy theory. This spectral sequence is called the kq-resolution and has signature
$E_{1}^{s,f,w}=\pi^{\mathbb{C}}_{s+f,w}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes f})\implies\pi_{s,w}^{\mathbb{C}}(\mathbb{S}_{\textup{kq}}^{\wedge}).$
Their analysis of this spectral sequence leads to, among many things, the $v_{1}$-torsion free component of $\pi^{\mathbb{C}}_{**}\mathbb{S}$. Toward computing the $E_{1}$-page, Culver–Quigley compute the ring of cooperations $\pi^{\mathbb{C}}_{**}(\textup{kq}\otimes\textup{kq})$ by the $\textup{H}\mathbb{F}_{2}\text{-}\textbf{mASS}(\text{kq}\otimes\text{kq})$ which takes the form
$E_{2}^{s,f,w}=\text{Ext}^{s,f,w}_{\mathcal{A}_{\mathbb{C}}^{\vee}}\left(\mathbb{M}_{2}^{\mathbb{C}},\textup{H}_{**}(\textup{kq}\otimes\textup{kq})\right)\implies\pi_{**}^{\mathbb{C}}(\textup{kq}\otimes\textup{kq}).$
@@ 1.2. Main results
The goal of this paper is to begin the study of the kq-resolution in $\mathbb{R}$-motivic homotopy theory. As a step in this direction, we compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$ as a module over $\pi_{**}^{\mathbb{R}}(\text{kq}).$
Theorem A (4.1, 8.7).
There is an isomorphism
$\pi_{**}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})\cong\bigoplus_{k\geq 0}\Sigma^{4k,2k}\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(\mathbb{M}_{2}^{\mathbb{R}},B_{0}^{\mathbb{R}}(k)\right),$
where $B_{0}^{\mathbb{R}}(k)$ denotes the $k$-th integral motivic Brown–Gitler comodule over $\mathbb{R}$ (see Section 2.3). We describe the right-hand completely, modulo $v_{1}$-torsion.
This result gives substantial control over the $n$-line of the $E_{1}$-page of the kq resolution; see Remark 8.8. The reader familiar with Mahowald’s analysis of the bo-resolution [Mah81] may be inclined to believe that this result may be proven in a similar fashion. This is not the case. Mahowald’s arguments make great use of integral Brown–Gitler spectra. However, there is currently no known construction of motivic integral Brown–Gitler spectra except in very particular cases (see Example 2.16, Remark 2.17), and so another method of analysis must be employed. As is the case over $\mathbb{C}$, we generally follow the strategy employed by [BOSS19], in which they compute the classical ring of cooperations for tmf using algebraic methods. We outline this strategy below.
@@
    We first introduce the integral motivic Brown–Gitler comodules $B_{0}^{\mathbb{R}}(k)$ and show that there are isomorphisms of $\mathcal{A}(1)^{\vee}_{\mathbb{R}}$-comodules:
@@
Steps (1) and (2) are direct consequences of [CQ21, Section 3]. Step (3) is the most difficult and occupies most of this paper. This is because the algebra $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})$ has more elaborate structure than its classical and $\mathbb{C}$-motivic counterparts. Indeed, when presenting the charts for our spectral sequence computations, we find it best to follow the strategy of Guillou–Hill–Isaksen–Ravenel [GHIR20] and organize each page by coweight $cw=stem-weight$. There is an element $\tau^{4}\in\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})$ which gives a periodicity allowing us to organize our data into 4 sets of charts. These computation are manageable with careful bookkeeping.
Through this process, we obtain the following result, where ksp denotes the very effective cover of $\Sigma^{4,2}\textup{KQ}$ (see Section 3):
Theorem B (Corollary 7.4).
There is an isomorphism:
$\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},B_{0}^{\mathbb{R}}(1))\cong\pi_{**}^{\mathbb{R}}(\textup{ksp}).$
This is an easy consequence of a stronger result which we show in a more general setting.
Theorem C (Proposition 3.7).
@@ 1.3. Future directions
This paper initiates the study of the kq-resolution in $\text{SH}(\mathbb{R})$. In particular, we calculate the $E_{1}$-page. While many of the differentials in this spectral sequence can be determined using the base change functor
$-\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C}),$
there is $\rho$-torsion which must be determined using the $C_{2}$-equivariant Betti realization functor
@@
along with information involving the $\text{ko}_{C_{2}}$-resolution. We plan to study this in future work. Additionally, the kq-resolution is intricately related to $v_{1}$-periodic homotopy. The work of Belmont–Isaksen–Kong [BIK24] determines a version of $v_{1}$-periodic homotopy theory in $\pi_{**}^{\mathbb{R}}\mathbb{S}$ by computing the homotopy groups $\pi_{**}^{\mathbb{R}}\textup{L}$, where L sits in a cofiber sequence
@@
It will be interesting to compare the homotopy groups of $\pi_{**}^{\mathbb{R}}\textup{L}$ with the $v_{1}$-periodic homotopy detected by the kq-resolution (see also Remark 3.8).
@@
In $C_{2}$-equivariant homotopy theory, the work of Li–Petersen–Tatum [LPT24] produces $C_{2}$-equivariant analogues of the classical integral Brown–Gitler spectra. These are spectra $\textup{H}\mathbb{Z}_{k}^{C_{2}}$ such that, as comodules over the the sub-Hopf algebra $\mathcal{A}(1)^{\vee}_{C_{2}}$ of the $C_{2}$-equivariant dual Steenrod algebra, there is an isomorphism
@@
where $B_{0}^{C_{2}}(k)$ is the $k$-th $C_{2}$-equivariant integral Brown–Gitler comodule. It was recently shown [LPT25] that there is a splitting of $\textup{ku}_{\mathbb{R}}$-module spectra:
@@
$E_{2}=\text{Ext}_{\mathcal{A}(1)^{\vee}_{C_{2}}}\left(\textup{H}_{\star}(\Sigma^{2\rho k}\textup{H}\mathbb{Z}_{k}^{C_{2}}),\textup{H}_{\star}(\textup{ko}_{C_{2}})\right)\implies[\Sigma^{2\rho k}\textup{ko}_{C_{2}}\otimes\textup{H}\mathbb{Z}_{k}^{C_{2}},\textup{ko}_{C_{2}}\otimes\textup{ko}_{C_{2}}]^{\textup{ko}_{C_{2}}}.$
It can be shown that the $E_{2}$-page of this spectral sequence decomposes into terms of the form
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{C_{2}}}(B_{0}^{C_{2}}(k),B_{0}^{C_{2}}(m)).$
This further splits as
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k),B_{0}^{\mathbb{R}}(m))\oplus\text{Ext}_{NC}(B_{0}^{C_{2}}(k),B_{0}^{C_{2}}(m)),$
where $\text{Ext}_{NC}$ refers to Ext over the negative cone. Our computations in this paper offer progress towards the $\mathbb{R}$-motivic Ext groups in this decomposition. These ideas will be further investigated in future work, joint with Li, Petersen, and Tatum.
1.3.1. The ring of cooperations over general bases
The work of Kylling [Kyl15] determines the group $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{F}_{3}}}(\mathbb{M}_{2}^{\mathbb{F}_{3}},\mathbb{M}_{2}^{\mathbb{F}_{3}})$. In forthcoming work, we use the methods developed in this paper to determine the ring of cooperations $\pi_{**}^{\mathbb{F}_{3}}(\textup{kq}\otimes\textup{kq})$. A remarkable result of Bachmann–Østvær [BØ22] gives a useful pullback square that allows us to make a deeper statement. Namely, if $\textup{E}\in\text{SH}(\mathbb{Z}[1/2])^{\text{cell}}$, then there is a pullback square
[…]
where $\text{map}_{F}(\mathbb{S},\textup{E})$ denotes the ordinary spectrum of maps between $\mathbb{S}$ and E in $\text{SH}(F)$ and and the morphisms are a combination of base change and Brauer liftings. The computation of $\pi_{**}^{F}(\textup{kq}\otimes\textup{kq})$ for $F=\mathbb{C},\mathbb{R},\mathbb{F}_{3}$ allows one to access the arithmetic invariant $\pi_{**}^{\mathbb{Z}[1/2]}(\textup{kq}\otimes\textup{kq})$. This bypasses the fact that, although we understand the dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{\mathbb{Z}[1/2]}$ (see [Spi18, Theorem 11.24], also [DØ23, Section 4.3]), there is currently no description of $\text{Ext}_{\mathcal{A}(1)_{\mathbb{Z}[1/2]}^{\vee}}(\mathbb{M}_{2}^{\mathbb{Z}[1/2]})$.

### v2

@@ Abstract.
Let kq denote the very effective cover of the motivic Hermitian K-theory spectrum. We analyze the ring of cooperations $\pi^{\mathbb{R}}_{**}(\textup{kq}\otimes\textup{kq})$ in the stable motivic homotopy category $\text{SH}(\mathbb{R})$, giving a full description in terms of Brown–Gitler comodules. To do this, we decompose the $\textup{E}_{2}$-page of the motivic Adams spectral sequence and show that it must collapse. The description of the $\textup{E}_{2}$-page is accomplished by a series of algebraic Atiyah–Hirzebruch spectral sequences which converge to the summands of the $\textup{E}_{2}$-page. Along the way, we prove a splitting result for the very effective symplectic K-theory ksp over any base field of characteristic not two.
@@ 1.1. Motivation
In stable homotopy theory, one of the most powerful tools we have for computing the homotopy groups of spheres is the Adams spectral sequence. For a ring spectrum E, the E-Adams spectral sequence (E-$\textbf{ASS}(\mathbb{S})$) has signature
$\textup{E}_{1}^{s,f}=\pi_{s+f}(\text{E}\otimes\overline{\text{E}}^{\otimes f})\implies\pi_{s}(\mathbb{S}_{\text{E}}^{\wedge}),$
where $\overline{\text{E}}$ is the cofiber of the unit map $\mathbb{S}\to\text{E}$. Under mild assumptions on E, this spectral sequence strongly converges to the homotopy groups of $\mathbb{S}_{\text{E}}^{\wedge}$, the E-nilpotent completion of the sphere spectrum [49]. To begin to compute with the E-$\textbf{ASS}(\mathbb{S})$, one must understand the smash powers $\text{E}^{\otimes f}$.
One may specialize from desiring to understand the entirety of $\pi_{*}(\mathbb{S})$ to desiring to understand a particular piece of $\pi_{*}(\mathbb{S})$. This is one of the payoffs of chromatic homotopy theory, which organizes the elements of the stable homotopy groups of spheres into $v_{n}$-periodic layers. The $v_{0}$-periodic elements of $\pi_{*}(\mathbb{S})$ are those which survive rationalization, hence the elements in degree 0. The next interesting layer of elements in the stable stems are those which are $v_{1}$-periodic, and so one may attempt to determine the $v_{1}$-torsion free component of $\pi_{*}(\mathbb{S})$ by some $\text{E-}\textbf{ASS}(\mathbb{S})$.
Let bo denote the connective cover of the real topological K-theory spectrum KO. This spectrum is well-enough behaved that one may begin to study its associated Adams spectral sequence. Mahowald and others have studied the bo-$\textbf{ASS}(\mathbb{S})$, known as the bo-resolution, at the prime 2 extensively [41, 36, 9]. Analysis of this spectral sequence has lead to, among many things, the $v_{1}$-torsion free component of $\pi_{*}(\mathbb{S})$. Key to Mahowald’s analysis is a spectrum level splitting of $\text{bo}\otimes\text{bo}$ as the following wedge sum:
The spectra $\textup{H}\mathbb{Z}_{k}^{cl}$ are finite complexes known as the integral Brown–Gitler spectra [14, 23].
@@
$\textup{E}_{1}^{s,f,w}=\pi^{F}_{s+f,w}(\textup{E}\otimes\overline{\textup{E}}^{\otimes f})\implies\pi_{s,w}^{F}(\mathbb{S}_{\textup{E}}^{\wedge}).$
[…]
$\textup{E}_{1}^{s,f,w}=\pi^{\mathbb{C}}_{s+f,w}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes f})\implies\pi_{s,w}^{\mathbb{C}}(\mathbb{S}_{\textup{kq}}^{\wedge}).$
Their analysis of this spectral sequence leads to, among many things, the $v_{1}$-torsion free component of $\pi^{\mathbb{C}}_{**}(\mathbb{S})$. Toward computing the $\textup{E}_{1}$-page, Culver–Quigley compute the ring of cooperations $\pi^{\mathbb{C}}_{**}(\textup{kq}\otimes\textup{kq})$ by the $\textup{H}\mathbb{F}_{2}\text{-}\textbf{mASS}^{\mathbb{C}}(\text{kq}\otimes\text{kq})$ which takes the form
$\textup{E}_{2}^{s,f,w}=\text{Ext}^{s,f,w}_{\mathcal{A}_{\mathbb{C}}^{\vee}}\left(\mathbb{M}_{2}^{\mathbb{C}},\textup{H}_{**}(\textup{kq}\otimes\textup{kq})\right)\implies\pi_{s,w}^{\mathbb{C}}(\textup{kq}\otimes\textup{kq}).$
@@ 1.2. Main results
The goal of this paper is to begin the study of the kq-resolution in $\mathbb{R}$-motivic homotopy theory. As a step in this direction, we compute the ring of cooperations $\pi_{**}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$ as a module over $\pi_{**}^{\mathbb{R}}(\text{kq}).$ Throughout the rest of this paper, let $\textbf{mASS}^{F}(X)$ denote the $F$-motivic $\text{H}\mathbb{F}_{2}$-based motivic Adams spectral sequence, where $\text{H}\mathbb{F}_{2}$ is the motivic Eilenberg-MacLane spectrum representing mod-2 motivic cohomology.
Theorem A (4.1, 8.4).
The $\textup{{mASS}}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$ has signature
$\textup{E}_{2}^{s,f,w}=\bigoplus_{k\geq 0}\Sigma^{4k,2k}\textup{Ext}_{\mathcal{A}(1)^{\vee}}^{s,f,w}(\mathbb{M}_{2}^{\mathbb{R}},B_{0}^{\mathbb{R}}(k))\implies\pi_{s,w}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$
and collapses on the $\textup{E}_{2}$-page. Here $B_{0}^{\mathbb{R}}(k)$ denotes the $k^{th}$ integral motivic Brown–Gitler comodule. We describe the $\textup{E}_{2}=\textup{E}_{\infty}$-page, modulo $v_{1}$-torsion, as a module over $\pi_{**}^{\mathbb{R}}\textup{(kq)}.$
The reader familiar with Mahowald’s analysis of the bo-resolution [41] may be inclined to believe that this result may be proven in a similar fashion. This is not the case. Mahowald’s arguments make great use of integral Brown–Gitler spectra. However, there is currently no known construction of motivic integral Brown–Gitler spectra except in very particular cases (see Example 2.8, Remark 2.9), and so another method of analysis must be employed. As is the case over $\mathbb{C}$, we generally follow the strategy employed by [10], in which they compute the classical ring of cooperations for tmf using algebraic methods. We outline this strategy below.
@@
    First, we introduce the integral motivic Brown–Gitler comodules $B_{0}^{\mathbb{R}}(k)$ and show that there are isomorphisms of $\mathcal{A}(1)^{\vee}_{\mathbb{R}}$-comodules:
@@
Steps (1) and (2) are direct consequences of [17, Section 3]. Step (3) is the most difficult and occupies most of this paper. This is because the algebra $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})$ has more elaborate structure than its classical and $\mathbb{C}$-motivic counterparts. Indeed, when presenting the charts for our spectral sequence computations, we find it best to follow the strategy of Guillou–Hill–Isaksen–Ravenel [24] and organize each page by coweight cw=stemweightcw=stem-weight. There is an element $\tau^{4}\in\text{Ext}^{0,0,-4}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})$ which gives a periodicity allowing us to organize our data into 4 sets of charts. These computation are manageable with careful bookkeeping.
As an application, we compute the $n$-line of the $\text{E}_{1}$-page of the kq-resolution over $\mathbb{R}$.
Theorem B (Proposition 8.6, 8.7).
The $\textup{{mASS}}^{\mathbb{R}}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n})$ has signature
$\textup{E}_{2}^{s,f,w}=\bigoplus_{K\in\mathcal{K}_{n}}\Sigma^{4|K|,2|K|}\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{s,f,w}(\mathbb{M}_{2}^{\mathbb{R}},B_{0}^{\mathbb{R}}(K))\implies\pi_{s,f}^{\mathbb{R}}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n}),$
and collapses on the $\textup{E}_{2}$-page, where $\mathcal{K}_{n}=\{K=(k_{1},\dots,k_{n}):k_{j}\geq 1\textup{ for all }j\}$, $|K|=\sum_{j=1}^{n}k_{j}$, and $B_{0}^{\mathbb{R}}(K)=\bigotimes_{j=1}^{n}B_{0}^{\mathbb{R}}(k_{j}).$
Our method of computation describes the $\text{E}_{1}$-page of the kq-resolution as a module over $\pi_{**}^{\mathbb{R}}(\text{kq})$ (reviewed in Section 6.1). We imagine that this structure will be useful in working with the kq-resolution in the style of [9].
Through our inductive procedure, we naturally obtain the following result, where ksp denotes the very effective cover of $\Sigma^{4,2}\textup{KQ}$ (see also Section 3).
Theorem C (Corollary 7.3).
The $\textup{{mASS}}^{\mathbb{R}}(\textup{ksp})$ has signature
$\textup{E}_{2}^{s,f,w}=\textup{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},B_{0}^{\mathbb{R}}(1))\implies\pi_{s,w}^{\mathbb{R}}(\textup{ksp}),$
and collapses on the $\textup{E}_{2}$-page.
This is an easy consequence of a splitting result which we show in a more general setting.
Theorem D (Proposition 3.3).
@@ 1.3. Future directions
This paper initiates the study of the kq-resolution in $\text{SH}(\mathbb{R})$. In particular, we calculate the $\textup{E}_{1}$-page. While many of the differentials in this spectral sequence can be determined using the base change functor
$\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C}),$
there is $\rho$-periodic data which must be determined using the $C_{2}$-equivariant Betti realization functor
@@
along with information involving the $\text{ko}_{C_{2}}$-resolution. We plan to study this in future work. Additionally, the kq-resolution is intricately related to $v_{1}$-periodic homotopy. The work of Belmont–Isaksen–Kong [11] determines a version of $v_{1}$-periodic homotopy theory in $\pi_{**}^{\mathbb{R}}\mathbb{S}$ by computing the homotopy groups $\pi_{**}^{\mathbb{R}}(\textup{L})$, where L sits in a cofiber sequence
@@
It will be interesting to compare the homotopy groups $\pi_{**}^{\mathbb{R}}(\textup{L})$ with the $v_{1}$-periodic homotopy detected by the kq-resolution (see also Remark 3.5).
@@
In $C_{2}$-equivariant homotopy theory, Li–Petersen–Tatum have produced $C_{2}$-equivariant analogues of the classical integral Brown–Gitler spectra [39]. These are finite spectra $\textup{H}\mathbb{Z}_{k}^{C_{2}}$ such that, as comodules over the the sub-Hopf algebra $\mathcal{A}(1)^{\vee}_{C_{2}}$ of the $C_{2}$-equivariant dual Steenrod algebra, there is an isomorphism
@@
where $B_{0}^{C_{2}}(k)$ is the $k^{th}$ $C_{2}$-equivariant integral Brown–Gitler comodule. It was recently shown [38] that there is a splitting of $\textup{ku}_{\mathbb{R}}$-module spectra
@@
$\textup{E}_{2}=\text{Ext}^{V,f}_{\mathcal{A}(1)^{\vee}_{C_{2}}}\left(\textup{H}_{\star}(\Sigma^{2\rho k}\textup{H}\mathbb{Z}_{k}^{C_{2}}),\textup{H}_{\star}(\textup{ko}_{C_{2}})\right)\implies[\Sigma^{2\rho k}\textup{ko}_{C_{2}}\otimes\textup{H}\mathbb{Z}_{k}^{C_{2}},\textup{ko}_{C_{2}}\otimes\textup{ko}_{C_{2}}]^{\textup{ko}_{C_{2}}}_{V}.$
Due to the isomorphism [24]
$\text{Ext}^{\star,*}_{\mathcal{A}(1)^{\vee}_{C_{2}}}(\mathbb{M}_{2}^{\text{C}_{2}},\mathbb{M}_{2}^{\text{C}_{2}})\cong\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{***}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})\oplus\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(NC,\mathbb{M}_{2}^{\mathbb{R}}),$
where $NC$ dentoes the “negative cone”, we expect that the Ext computations in this paper will be valuable in solving this problem. These ideas will be further investigated in future work, joint with Li, Petersen, and Tatum.
The ring of cooperations over general bases
One can also ask about the ring of cooperations over more general base schemes. A necessary input to using the methods of Section 4 to compute $\pi_{**}(\text{kq}\otimes\text{kq})$ is a computation of the algebra $\text{Ext}_{\mathcal{A}(1)^{\vee}}^{***}(\mathbb{M}_{2},\mathbb{M}_{2})$. In [34], Kylling determined $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{F}_{q}}}(\mathbb{M}_{2}^{\mathbb{F}_{q}},\mathbb{M}_{2}^{\mathbb{F}_{q}})$ for $\text{char}(\mathbb{F}_{q})\neq 2$, which allowed us to follow the program layed out in this paper and compute the ring of cooperations $\pi_{**}^{\mathbb{F}_{q}}(\textup{kq}\otimes\textup{kq})$ [46]. In forthcoming work with Petersen and Tatum, we compute both $\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}^{***}(\mathbb{M}_{2}^{F},\mathbb{M}_{2}^{F})$ and $\pi_{**}^{F}(\text{kq}\otimes\text{kq})$ when $F=\mathbb{Q}$ and $F=\mathbb{Q}_{p}$ for $p$ any prime [44].
The case of finite fields is particularly interesting when paired with our computation and the existing $\mathbb{C}$-motivic computation. A remarkable result of Bachmann–Østvær [5] gives a useful pullback square that allows us to make a deeper statement. Namely, if $\textup{E}\in\text{SH}(\mathbb{Z}[1/2])^{\text{cell}}$, then there is a pullback square
$\text{map}_{\mathbb{Z}[1/2]}(\mathbb{S},\textup{E})_{2}^{\wedge}$$\text{map}_{\mathbb{R}}(\mathbb{S},\textup{E})_{2}^{\wedge}$$\text{map}_{\mathbb{F}_{3}}(\mathbb{S},\textup{E})_{2}^{\wedge}$$\text{map}_{\mathbb{C}}(\mathbb{S},\textup{E})_{2}^{\wedge}$$\scriptstyle{\lrcorner}$
where $\text{map}_{F}(\mathbb{S},\textup{E})$ denotes the (ordinary) spectrum of maps between $\mathbb{S}$ and E in $\text{SH}(F)$. The computation of $\pi_{**}^{F}(\textup{kq}\otimes\textup{kq})$ for $F=\mathbb{C},\mathbb{R},\mathbb{F}_{3}$ combined with this square gives a long exact sequence which allows one to access the arithmetic invariant $\pi_{**}^{\mathbb{Z}[1/2]}(\textup{kq}\otimes\textup{kq})$. This bypasses the fact that, although we understand the dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{\mathbb{Z}[1/2]}$ (see [54, Theorem 11.24], also [20, Section 4.3]), there is currently no description of $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{Z}[1/2]}^{\vee}}(\mathbb{M}_{2}^{\mathbb{Z}[1/2]},\mathbb{M}_{2}^{\mathbb{Z}[1/2]})$.